OpenStudy (henrietepurina):

Algebra 1 question :D Medal + Fan!

OpenStudy (henrietepurina):

The function H(t) = -16t^2 + vt + s shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Question: What is the maximum height that the projectile will reach? Show your work.

OpenStudy (henrietepurina):

@AlexandervonHumboldt2 @bibby

OpenStudy (henrietepurina):

@paki

OpenStudy (anonymous):

you have no numbers to work with here so your answer must be in terms of \(v\) and \(s\)

OpenStudy (henrietepurina):

ok will do...

OpenStudy (anonymous):

which is kind of annoying, but we can do it

OpenStudy (henrietepurina):

what shall I begin with?

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[\frac{v}{32}\]

OpenStudy (henrietepurina):

ok...

OpenStudy (anonymous):

second coordinate (the max you are looking for) is what you get when you replace \(t\) by \(\frac{v]{32}\)

OpenStudy (anonymous):

oops replace \(t\) by \(\frac{v}{32}\)

OpenStudy (henrietepurina):

nah thats ok..... hmm, now how would you replace t by (v/32)?

OpenStudy (henrietepurina):

(t/32)?

OpenStudy (anonymous):

no just plug it in

OpenStudy (henrietepurina):

no wait.... that would just literally mean replace

OpenStudy (henrietepurina):

yeah, yeah... i'm a wee bit slow

OpenStudy (henrietepurina):

s+((v^2)/(64))

OpenStudy (henrietepurina):

this is point 2, not max height right?

OpenStudy (anonymous):

that is the max height yes

OpenStudy (anonymous):

depends on both s and v

OpenStudy (henrietepurina):

thanks how would I show my work?

OpenStudy (henrietepurina):

Since first coordinate of the vertex is always −((b)/(2a)) which in my case is ((v)(32)). To find the max, you have to replace t by ((v)(32)). Then you just plug t in, to get s+((v^2)/(64)). Thus, the max height is s+((v^2)/(64))

OpenStudy (anonymous):

looks good to me

OpenStudy (henrietepurina):

Thanks... can you live through 2 more questions?

OpenStudy (anonymous):

lol maybe lets see

OpenStudy (henrietepurina):

Both of these have to do with the original statement. Part C: Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know?

OpenStudy (anonymous):

\[g(t)=31+32.2t\]?

OpenStudy (henrietepurina):

Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in feet, of the object from the ground at time t seconds. That's all I got!

OpenStudy (anonymous):

ok then i think we are missing something the first line about H has this The function H(t) = -16t^2 + vt + s shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second.

OpenStudy (anonymous):

it has no numbers in it at all

OpenStudy (anonymous):

so without knowing \(v\) or \(s\) we can't possibly do this problem we can't see where they intersect i am thinking somewhere along the line you were told v and s, and we missed it somehow

OpenStudy (anonymous):

if we do know v and s we have to change the answer to the first question too

OpenStudy (henrietepurina):

Yeah.... but remember this is part c and d of the question we just answered.

OpenStudy (henrietepurina):

The projectile was launched from a height of 96 feet with an initial velocity of 80 feet per second.

OpenStudy (henrietepurina):

FOUND IT :D

OpenStudy (anonymous):

oh lord have mercy ok lets start at the top (quickly)

OpenStudy (henrietepurina):

ok then

OpenStudy (henrietepurina):

haha... "oh lord have mercy"

OpenStudy (anonymous):

forget the first answer with the s and v, now we do it with numbers \[h(t)=-16t^2+80t+96\] first coordianate is \[-\frac{b}{2a}\]which is now \[\frac{80}{32}=2.5\]

OpenStudy (anonymous):

second coordinate, which is the maximum, is \(h(2.5)\)which i will let you compute (use a calculator )

OpenStudy (henrietepurina):

so 2.5 = −16t^2+80t+96 or 16*2.5^2+80*2.5+96

OpenStudy (anonymous):

\[g(t)=31+32.2t\] they will intersect when \(h(t)=g(t)\) so we set them equal and solve

OpenStudy (anonymous):

you are missing a minus sign \[-16*(2.5)^2+80*2.5+96 \]

OpenStudy (anonymous):

i get 196 http://www.wolframalpha.com/input/?i=-16*2.5^2%2B80*2.5%2B96

OpenStudy (henrietepurina):

ok then so for the first one we did.... the answer is 196?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

max height is 196 when \(t=2.5\)

OpenStudy (henrietepurina):

Mind helping me put this answer in to a few sentences?

OpenStudy (anonymous):

say exactly what you said before, but use the numbers rather than the s and t

OpenStudy (henrietepurina):

oh ok... gimme a sec then :P

OpenStudy (anonymous):

Since first coordinate of the vertex is always −((b)/(2a)) which in my case is 80/32=2.5. To find the max, you have to replace t by 2.5. Then you just plug t in, to get -16(2.5)^2+80*2.5=96=196 Thus, the max height is 196 feet

OpenStudy (anonymous):

ok now for C

OpenStudy (henrietepurina):

Ignore my reply, that was for the first answer :P

OpenStudy (anonymous):

set \[h(t)=-16t^2+80t+96=g(t)=31+32.2t\] and solve for \(t\)

OpenStudy (anonymous):

it says use a table, we will use technology

OpenStudy (henrietepurina):

haha... ok then :D

OpenStudy (anonymous):

we graph them both we see that they do intersect, somewhere around \(t=4\) i.e. in 4 seconds

OpenStudy (henrietepurina):

How would you find the approximate solution to the equation H(t) = g(t)? −16t^2+80t+96=31+32.2t

OpenStudy (anonymous):

set them equal and ask wolfram

OpenStudy (henrietepurina):

ok

OpenStudy (anonymous):

Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know? since they meet around 4 second, the projectile is going down that is because it reaches it's maximum height at 2.5 seconds and starts going down after

OpenStudy (henrietepurina):

t=−1.0149925261473127,4.0024925261473125 thats what I got...

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

did you check the link i sent ?

OpenStudy (henrietepurina):

yeah, but I entered it incorrectly :P

OpenStudy (henrietepurina):

so they intersect at 2?

OpenStudy (henrietepurina):

x = 2?